OpenStudy (anonymous):

plz prove log(1+x)/(1-x) an odd function

OpenStudy (anonymous):

ok first off \[\ln(\frac{(1+x)}{(1-x)})=\ln(1+x)-\ln(1-x)\]\][

OpenStudy (anonymous):

now replace x by -x and you will have it \[f(x)=\ln(1+x)-\ln(1-x)\] \[f(-x)=\ln(1-x)-\ln(1+x)=-(\ln(1+x)-\ln(1-x))=-f(x)\]